OpenStudy (anonymous):

a carpenter cut four lengths of wood the lengths were 36 3/4 in 36 3/8 in 37 1/2 in and z in if the mean of the lengths is 36 5/8 in what was the length of the fourth piece of wood the carpenter cut answers only

OpenStudy (anonymous):

i hate mixed numbers. use decimals and a calculator you have \[\frac{a+b+c+z}{4}=36\tfrac{5}{8}\] so \[a+b+c+z=4\times 36\tfrac{5}{8}\] you know the numbers \(a,b,c\) so solve for \(z\)

OpenStudy (anonymous):

I really dont understand this please help

OpenStudy (anonymous):

the "mean" is what you get when you add up and divide by the number of numbers you have. in every day english the "average"

OpenStudy (anonymous):

you only know 3 out of the 4 numbers, the 4th is \(z\) and that is the number you need to solve for you are given the mean is \(36\tfrac{5}{8}\) so you know \[\frac{36\tfrac{3}{4}+36\tfrac{3}{8}+37\tfrac{1}{2}+z}{4}=36\tfrac{5}{8}\]

OpenStudy (anonymous):

multiply both sizes by 4 to get \[36\tfrac{3}{4}+36\tfrac{3}{8}+37\tfrac{1}{2}+z=4\times 36\tfrac{5}{8}\]

OpenStudy (anonymous):

add up the mess on the left, subtract that number from the right, and you will have \(z\)

OpenStudy (anonymous):

and the answer is????????